OpenStudy (anonymous):

The sum of 4 times a number and 14 is 16.What is the number?

OpenStudy (anonymous):

It's just 4x+14=16 solve for x

OpenStudy (inowalst):

@Daina900 Can you solve what @iambatman just told you?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Show an attempt, doing algebra you do the opposite operation of what already exists to both sides to get the variable x on its on.

OpenStudy (inowalst):

4x+14=16 Do you know what to do first? Try at least.

OpenStudy (anonymous):

4x+14=16 4x=16-14 4x=2 x=2/4 x=0.5